# Language Model with Retrieval-Augmented Generation

### v1

@@ 1 Introduction
In recent years, embedding models have made remarkable strides, becoming foundational components for a wide range of natural language processing tasks(Su et al., 2023) (Xiao et al., 2024) (Wang et al., 2023). These models excel at capturing semantic relationships within text data, enabling robust retrieval and matching capabilities. The success of text embeddings has naturally extended to explorations in multimodal domains, leading to significant advancements in text-image embedding models that effectively align semantic spaces across modalities (Chen et al., 2022) (Yasunaga et al., 2023) (Hu et al., 2023). However, despite the pervasive presence of speech content in real-world applications, the integration of speech modalities into embedding-based retrieval systems remains underexplored, highlighting a critical gap in multimodal information processing.
@@
This challenge becomes particularly pronounced in speech-based retrieval-augmented generation (RAG) systems, which have emerged as a promising approach to augment speech large language models (SLLMs) with external knowledge (Jacqmin et al., 2023). Current speech RAG systems predominantly adopt a two-stage paradigm: automatic speech recognition (ASR) followed by text-based RAG as shown in Fig 1. While this approach leverages well-established components, it suffers from two major limitations. First, the sequential processing of ASR and text-based RAG significantly increases system latency, rendering it impractical for real-time applications where rapid responses are critical. Second, and more fundamentally, this cascaded architecture is prone to error propagation. ASR errors—particularly in challenging acoustic environments or with non-standard speech—inevitably degrade downstream retrieval performance. This cascading effect can result in entirely irrelevant document retrieval, even when ASR errors impact only a few key terms.
Several attempts have been made to address these challenges through intermediate solutions, such as optimizing ASR for retrieval-specific metrics or implementing early fusion techniques (Serdyuk et al., 2018) (Szymański et al., 2023). However, these approaches still maintain the fundamental separation between speech processing and document retrieval, failing to capture the direct relationship between acoustic patterns and document semantics. Furthermore, existing methods often struggle with the inherent temporal nature of speech and the challenge of aligning variable-length speech sequences with document embeddings.
@@ 2.1 Multimodal Large Language Models
Recent advances in multimodal large language models (MLLMs) (Li et al., 2022) (Li et al., 2023) (Liu et al., 2024) have demonstrated impressive capabilities in processing visual and linguistic information. Models like LLaVA (Liu et al., 2024) have established effective architectures combining LLMs with visual encoders through carefully designed projection layers. This architecture typically involves a two-stage training process: first aligning visual-textual representations, then fine-tuning on instruction data. While most efforts have focused on vision-language models, some recent work has explored speech-language models by adapting similar architectures (Zhang et al., 2023) (Fang et al., 2024). However, these models primarily focus on speech understanding and generation rather than creating robust speech embeddings for retrieval tasks. Our work bridges this gap by introducing an embedding-focused architecture specifically designed for speech-text alignment.
@@ 2.2 Multimodal Embeddings
In the field of multimodal embeddings, CLIP (Radford et al., 2021) has established a strong foundation for vision-language understanding through contrastive learning on large-scale image-text pairs. Its success has inspired numerous follow-up works in visual-textual alignment and retrieval. While vision-language embedding models have made significant progress, the speech domain remains relatively unexplored. Current speech-text retrieval systems typically rely on a cascaded approach, first converting speech to transcripts through ASR Hinton et al. (2012); Prabhavalkar et al. (2024); Zhang et al. (2024) before applying text-based retrieval methods. This two-stage process, while straightforward, introduces latency and error propagation issues. The lack of end-to-end speech embedding models, analogous to CLIP’s role in vision, represents a significant gap in multimodal understanding. Our work addresses this limitation by introducing a unified speech embedding framework that directly aligns speech with textual content, demonstrating that end-to-end speech-text embeddings can achieve superior performance while reducing computational overhead.
@@ 3.1.1 Speech-Text Alignment Pre-Training
To bridge the modality gap between speech and text features, we introduce an adaptation module inspired by LLaVA’s (Liu et al., 2024) design. This module consists of two components: a temporal convolutional layer for dimension reduction and a Multi-Layer Perceptron (MLP) for feature projection. The adaptation process can be formulated as:
@@
where $T^{\prime}$ is the reduced sequence length after convolution, and the output dimension matches BERT’s(Devlin, 2018) hidden size $d_{t}$. The Conv1D layer uses a kernel size of k and stride s to effectively capture local temporal dependencies while reducing the sequence length. The MLP consists of two linear layers with a GELU activation function:
@@
2:  $h_{s}\leftarrow f_{s}(x_{s})$
3:  $z_{t}\leftarrow f_{t}(x_{t})$
4:  $h_{s}^{\prime}\leftarrow\text{Conv1D}(h_{s})$
5:  $z_{s}\leftarrow\text{MLP}(h_{s}^{\prime})$
6:  $\mathcal{L}_{\text{pre}}\leftarrow\big{\|}\frac{1}{T^{\prime}}\sum_{i=1}^{T^{\prime}}z_{s}^{i}-\frac{1}{L}\sum_{i=1}^{L}z_{t}^{i}\big{\|}_{2}^{2}$
@@
8:  $q\leftarrow\text{Linear}(\text{Adapter}(f_{s}(x_{s})))$
9:  $k^{+}\leftarrow\text{Linear}(f_{t}(d^{+}))$
10:  for $i\leftarrow 1$ to $N$ do
@@ 3.1.2 Speech-Text Retrieval Fine-Tuning
In the fine-tuning stage, we employ multi-task hybrid loss to optimize the alignment between speech queries and relevant document embeddings. For an speech query $x_{s}$ and its corresponding document $d^{+}$, along with a set of negative documents $\{d^{-}_{i}\}_{i=1}^{N}$, we first encode them using our pre-trained encoders. Inspired by text-embedding-v3 (OpenAI, 2024) and Piccolo2 (Huang et al., 2024), we scale up the embedding vector dimension to enhance model capacity. A shared linear layer is then added to the final layers of both the speech and text encoders. The operation can be formulated as:
@@
For retrieval and reranking tasks, we use InfoNCE loss (Gutmann and Hyvärinen, 2010) with cosine similarity:
@@
=$\mathcal{L}=\begin{cases}\mathcal{L}_{\text{cls}}&\text{if task is classification or clustering}\\ \mathcal{L}_{\text{sts}}&\text{if task is STS or pair-classification}\\ \mathcal{L}_{\text{re}}&\text{if task is retrieval or reranking}\end{cases}$ (12)
@@ 3.2.2 Speech-Text Alignment with CTC Loss
We also explore using Connectionist Temporal Classification (CTC) loss (Graves et al., 2006) for temporal alignment between speech and text sequences:
@@ 4.1 Implementation Details
For text embedding, we utilize piccolo-large-zh-v2 (Huang et al., 2024) as our text encoder, which provides 1024-dimensional text representations. For speech encoding, we employ Whisper-large-v3 encoder (Radford et al., 2023), leveraging its robust speech feature extraction capabilities trained on large-scale speech-text data.
@@
We employ the AdamW (Loshchilov and Hutter, ) optimizer with weight decay of 0.01 and a linear learning rate scheduler with 10% warmup steps. The temperature parameter $\tau$ in contrastive learning is set to 0.07.
@@ 4.2 Datasets
In the training process, different datasets are used at each stage. For the first stage, which is the alignment phase, we collected publicly available data (Emilia) (He et al., 2024) and a subset of data from the internet, totaling 170k hours. In the second stage, we used text training data from PiccoloV2, which includes both open-source data and data generated using specific methods. For each query, we generated data using CosyVoice-300M (Du et al., 2024). During the generation process, to ensure the model’s generalization ability, we also used six different voice tones for generation. In the second stage, a total of 80k hours of data were used for contrastive learning training. Inspired by Emilia, we also built a data filtering pipeline. In the first stage of data filtering, we applied several techniques, including Source Separation, Speaker Diarization, and Fine-grained Segmentation by VAD, to split the speech into clean speech from a specific speaker. Finally, we used ASR to obtain the corresponding text. For both the first and second stages of data, we employed DNSMOS P.835 OVRL(Reddy et al., 2022), keeping only the speech with scores above 3.0. Additionally, any speech segments with average phoneme durations exceeding 1.5 times the interquartile range (IQR) were discarded.
@@ 4.3.1 CMTEB
As an authoritative benchmark for evaluating text embedding tasks, MTEB (Massive Text Embedding Benchmark) (Muennighoff et al., 2022) was introduced by Muennighoff et al. in 2022. For Chinese scenarios, Xiao et al. developed CMTEB (Chinese Massive Text Embedding Benchmark) (Xiao et al., 2024) in 2023, which contains 35 datasets covering 6 major categories: Classification, Clustering, Pair Classification, Rerank, Retrieval, and Semantic Textual Similarity (STS). In our experiments, we utilized CosyVoice text-to-speech technology to convert these textual data into corresponding speech data for evaluating our model’s performance. We show the results in Table 1. As shown in the table, our method outperforms the two-stage ASR approach, regardless of whether Piccolo or the current state-of-the-art method, Conan (Li et al., 2024), is used for alignment. We believe that if the alignment text model uses Conan, the performance is likely to improve further.
@@ 4.4 Ablation Study
Additionally, in the experiments focusing solely on Stage 2 training, we compared three pretraining models: Hubert (Hsu et al., 2021), SenseVoice-small (FunAudioLLM, 2024), and Whisper (Radford et al., 2023). The results showed that Whisper outperformed the other two significantly, likely due to its use of a large amount of speech data during pretraining.

### v2

@@ 1 Introduction
In recent years, embedding models have made remarkable strides, becoming foundational components for a wide range of natural language processing tasks(su2023one) (cpack) (wang2023improving). These models excel at capturing semantic relationships within text data, enabling robust retrieval and matching capabilities. The success of text embeddings has naturally extended to explorations in multimodal domains, leading to significant advancements in text-image embedding models that effectively align semantic spaces across modalities (chen2022murag) (yasunaga2023retrieval) (hu2023reveal). However, despite the pervasive presence of speech content in real-world applications, the integration of speech modalities into embedding-based retrieval systems remains underexplored, highlighting a critical gap in multimodal information processing.
@@
This challenge becomes particularly pronounced in speech-based retrieval-augmented generation (RAG) systems, which have emerged as a promising approach to augment speech large language models (SLLMs) with external knowledge (jacqmin2023olisia). Current speech RAG systems predominantly adopt a two-stage paradigm: automatic speech recognition (ASR) followed by text-based RAG as shown in Fig 1. While this approach leverages well-established components, it suffers from two major limitations. First, the sequential processing of ASR and text-based RAG significantly increases system latency, rendering it impractical for real-time applications where rapid responses are critical. Second, and more fundamentally, this cascaded architecture is prone to error propagation. ASR errors—particularly in challenging acoustic environments or with non-standard speech—inevitably degrade downstream retrieval performance. This cascading effect can result in entirely irrelevant document retrieval, even when ASR errors impact only a few key terms.
Several attempts have been made to address these challenges through intermediate solutions, such as optimizing ASR for retrieval-specific metrics or implementing early fusion techniques (serdyuk2018towards) (szymanski2023aren). However, these approaches still maintain the fundamental separation between speech processing and document retrieval, failing to capture the direct relationship between acoustic patterns and document semantics. Furthermore, existing methods often struggle with the inherent temporal nature of speech and the challenge of aligning variable-length speech sequences with document embeddings.
@@ 2.1 Multimodal Large Language Models
Recent advances in multimodal large language models (MLLMs) (li2022blip) (li2023blip) (liu2024visual) have demonstrated impressive capabilities in processing visual and linguistic information. Models like LLaVA (liu2024visual) have established effective architectures combining LLMs with visual encoders through carefully designed projection layers. This architecture typically involves a two-stage training process: first aligning visual-textual representations, then fine-tuning on instruction data. While most efforts have focused on vision-language models, some recent work has explored speech-language models by adapting similar architectures (zhang2023speechgpt) (fang2024llama). However, these models primarily focus on speech understanding and generation rather than creating robust speech embeddings for retrieval tasks. Our work bridges this gap by introducing an embedding-focused architecture specifically designed for speech-text alignment.
@@ 2.2 Multimodal Embeddings
In the field of multimodal embeddings, CLIP (radford2021learning) has established a strong foundation for vision-language understanding through contrastive learning on large-scale image-text pairs. Its success has inspired numerous follow-up works in visual-textual alignment and retrieval. While vision-language embedding models have made significant progress, the speech domain remains relatively unexplored. Current speech-text retrieval systems typically rely on a cascaded approach, first converting speech to transcripts through ASR Hinton2012; Prabhavalkar2023; DBLP:conf/icassp/ZhangZZZL24 before applying text-based retrieval methods. This two-stage process, while straightforward, introduces latency and error propagation issues. The lack of end-to-end speech embedding models, analogous to CLIP’s role in vision, represents a significant gap in multimodal understanding. Our work addresses this limitation by introducing a unified speech embedding framework that directly aligns speech with textual content, demonstrating that end-to-end speech-text embeddings can achieve superior performance while reducing computational overhead.
@@ 3.1.1 Speech-Text Alignment Pre-Training
To bridge the modality gap between speech and text features, we introduce an adaptation module inspired by LLaVA’s (liu2024visual) design. This module consists of two components: a temporal convolutional layer for dimension reduction and a Multi-Layer Perceptron (MLP) for feature projection. The adaptation process can be formulated as:
@@
where $T^{\prime}$ is the reduced sequence length after convolution, and the output dimension matches BERT’s(devlin2018bert) hidden size $d_{t}$. The Conv1D layer uses a kernel size of k and stride s to effectively capture local temporal dependencies while reducing the sequence length. The MLP consists of two linear layers with a GELU activation function:
@@
[…]
@@
[…]
10:for $i\leftarrow 1$ to NN do
@@ 3.1.2 Speech-Text Retrieval Fine-Tuning
In the fine-tuning stage, we employ multi-task hybrid loss to optimize the alignment between speech queries and relevant document embeddings. For an speech query $x_{s}$ and its corresponding document $d^{+}$, along with a set of negative documents $\{d^{-}_{i}\}_{i=1}^{N}$, we first encode them using our pre-trained encoders. Inspired by text-embedding-v3 (openai2024embedding) and Piccolo2 (huang2024piccolo2), we scale up the embedding vector dimension to enhance model capacity. A shared linear layer is then added to the final layers of both the speech and text encoders. The operation can be formulated as:
@@
For retrieval and reranking tasks, we use InfoNCE loss (gutmann2010noise) with cosine similarity:
@@
[…]
@@ 3.2.2 Speech-Text Alignment with CTC Loss
We also explore using Connectionist Temporal Classification (CTC) loss (graves2006connectionist) for temporal alignment between speech and text sequences:
@@ 4.1 Implementation Details
For text embedding, we utilize piccolo-large-zh-v2 (huang2024piccolo2) as our text encoder, which provides 1024-dimensional text representations. For speech encoding, we employ Whisper-large-v3 encoder (radford2023robust), leveraging its robust speech feature extraction capabilities trained on large-scale speech-text data.
@@
We employ the AdamW (loshchilovdecoupled) optimizer with weight decay of 0.01 and a linear learning rate scheduler with 10% warmup steps. The temperature parameter $\tau$ in contrastive learning is set to 0.07.
@@ 4.2 Datasets
In the training process, different datasets are used at each stage. For the first stage, which is the alignment phase, we collected publicly available data (Emilia) (he2024emilia) and a subset of data from the internet, totaling 170k hours. In the second stage, we used text training data from PiccoloV2, which includes both open-source data and data generated using specific methods. For each query, we generated data using CosyVoice-300M (du2024cosyvoice). During the generation process, to ensure the model’s generalization ability, we also used six different voice tones for generation. In the second stage, a total of 80k hours of data were used for contrastive learning training. Inspired by Emilia, we also built a data filtering pipeline. In the first stage of data filtering, we applied several techniques, including Source Separation, Speaker Diarization, and Fine-grained Segmentation by VAD, to split the speech into clean speech from a specific speaker. Finally, we used ASR to obtain the corresponding text. For both the first and second stages of data, we employed DNSMOS P.835 OVRL(reddy2022dnsmos), keeping only the speech with scores above 3.0. Additionally, any speech segments with average phoneme durations exceeding 1.5 times the interquartile range (IQR) were discarded.
@@ 4.3.1 CMTEB
As an authoritative benchmark for evaluating text embedding tasks, MTEB (Massive Text Embedding Benchmark) (muennighoff2022mteb) was introduced by Muennighoff et al. in 2022. For Chinese scenarios, Xiao et al. developed CMTEB (Chinese Massive Text Embedding Benchmark) (cpack) in 2023, which contains 35 datasets covering 6 major categories: Classification, Clustering, Pair Classification, Rerank, Retrieval, and Semantic Textual Similarity (STS). In our experiments, we utilized CosyVoice text-to-speech technology to convert these textual data into corresponding speech data for evaluating our model’s performance. We show the results in Table 1. As shown in the table, our method outperforms the two-stage ASR approach, regardless of whether Piccolo or the current state-of-the-art method, Conan (li2024conan), is used for alignment. We believe that if the alignment text model uses Conan, the performance is likely to improve further.
@@ 4.4 Ablation Study
Additionally, in the experiments focusing solely on Stage 2 training, we compared three pretraining models: Hubert (hsu2021hubert), SenseVoice-small (sensevoice2024), and Whisper (radford2023robust). The results showed that Whisper outperformed the other two significantly, likely due to its use of a large amount of speech data during pretraining.